### v1

@@ 4.1.2 Model Selection and Training Regimen
We conducted experiments on both English and Spanish datasets. The specific models used for each language are detailed in Tables 4 and 5. These models were selected based on their suitability and previous success in addressing similar NLP challenges as addressed in Section 2.1. All models were fine-tuned using the AdamW optimizer with a learning rate of 3e-5 and a weight decay of 0.01. We trained for a maximum of 15 epochs per fold, employing early stopping with a patience of 3 epochs based on the performance on a held-out development set. The maximum sequence length was set to 100 tokens. A batch size of 32 was used for all experiments. The choice of these hyperparameters was guided by a coarse grid search on a separate development set, aiming to optimize performance while preventing overfitting.
Table 4: Pre-trained Transformer Models Used for English.
@@
Table 5: Pre-trained Transformer Models Used for Spanish.
@@
Through this fine-tuning exercise, we aim to establish strong, reproducible baselines for the PolyHope V2. The results obtained will quantify the effectiveness of Transformer models for hope speech detection and, critically, sarcasm identification. This comparative analysis will provide a benchmark against which to evaluate the performance gains achieved by our proposed ZSL and FSL methodologies leveraging LLMs.
@@ 4.2 Zero-Shot and Few-Shot Learning with Large Language Models
We selected GPT-4 [43, 44] and Llama 3 [45] for their superior performance in various natural language understanding and generation tasks. GPT-4, known for its advanced reasoning and generalization abilities, serves as a high-performing benchmark. Llama 3, a powerful open-source LLM, offers an alternative perspective and allows for greater control and customization.
@@ 4.2.2 Implementation Details
    messages=[{"role": "system", "content": "You are an NLP model trained
    for text classification."},
@@
By evaluating GPT-4 and Llama 3 in ZSL and FSL settings, we aim to assess the potential of LLMs for hope speech and sarcasm detection without task-specific training. For the FSL setting, we have taken a sample of 10 per label in our prompt draft and predicted on the rest of the datasets for both languages. The results will provide insights into the strengths and limitations of these models and inform future research directions in this area. We expect that FSL will outperform ZSL, and that GPT-4 and Llama 3 will exhibit different strengths and weaknesses due to their architectural differences and training data. A comparison with the fine-tuned Transformer baselines will reveal the trade-offs between these two approaches and highlight the potential for combining them to achieve even better performance.
5 Results
Table 6: Results for English language by fine tuning PLMs.
Model Avg. Weighted Scores Avg. Macro Scores Accuracy
Precision Recall F1-score Precision Recall F1-score
Binary Hope Speech Detection
RoBERTa 0.8656 0.8650 0.8651 0.8643 0.8653 0.8646 0.8650
Albert 0.8330 0.8327 0.8326 0.8326 0.8314 0.8317 0.8327
Electra 0.8530 0.8525 0.8526 0.8517 0.8526 0.8520 0.8525
DistilBERT 0.8604 0.8599 0.8599 0.8591 0.8601 0.8594 0.8599
Multiclass Hope Speech Detection
RoBERTa 0.7929 0.7885 0.7900 0.7551 0.7645 0.7587 0.7885
Albert 0.7728 0.7741 0.7723 0.7474 0.7251 0.7347 0.7741
Electra 0.7705 0.7626 0.7657 0.7217 0.7361 0.7277 0.7626
DistilBERT 0.7837 0.7813 0.7822 0.7480 0.7450 0.7461 0.7813
Table 7: Results for English language using prompt-learning.
Model Approach Avg. Weighted Scores Avg. Macro Scores Accuracy
Precision Recall F1-score Precision Recall F1-score
Binary Hope Speech Detection
Llama3 FSL 0.7626 0.7401 0.7379 0.7566 0.7481 0.7391 0.7401
ZSL 0.6743 0.6695 0.6678 0.6728 0.6592 0.6603 0.6784
GPT-4 FSL 0.7932 0.7789 0.7795 0.7861 0.7724 0.7711 0.7823
ZSL 0.7717 0.7706 0.7692 0.7724 0.7662 0.7674 0.7706
Multiclass Hope Speech Detection
Llama3 FSL 0.5090 0.4374 0.4244 0.4677 0.3933 0.3494 0.4374
ZSL 0.3886 0.3741 0.3698 0.3517 0.3086 0.2979 0.3772
GPT-4 FSL 0.4872 0.4594 0.4485 0.4388 0.4096 0.3963 0.4481
ZSL 0.5445 0.4354 0.4282 0.4638 0.4497 0.3690 0.4354
Table 8: Results for Spanish language by fine tuning PLMs.
Model Avg. Weighted Scores Avg. Macro Scores Accuracy
Precision Recall F1-score Precision Recall F1-score
Binary Hope Speech Detection
Roberta 0.8410 0.8401 0.8401 0.8403 0.8400 0.8397 0.8401
Albert 0.8426 0.8425 0.8425 0.8420 0.8422 0.8420 0.8425
Electra 0.8332 0.8308 0.8309 0.8314 0.8323 0.8308 0.8308
DistilBERT 0.7699 0.7697 0.7697 0.7691 0.7692 0.7690 0.7697
Multiclass Hope Speech Detection
Roberta 0.7603 0.7604 0.7593 0.7343 0.7242 0.7281 0.7604
Albert 0.7587 0.7585 0.7581 0.7306 0.7229 0.7261 0.7585
Electra 0.7493 0.7410 0.7444 0.7038 0.7176 0.7097 0.7410
DistilBERT 0.6501 0.6530 0.6509 0.6154 0.5986 0.6061 0.6530
Table 9: Results for Spanish language using prompt-learning (estimated values filled).
Model Approach Avg. Weighted Scores Avg. Macro Scores Accuracy
Precision Recall F1-score Precision Recall F1-score
Binary Hope Speech Detection
Llama3 FSL 0.7415 0.7221 0.7194 0.7374 0.7283 0.7205 0.7221
ZSL 0.6627 0.6472 0.6431 0.6588 0.6410 0.6359 0.6489
GPT-4 FSL 0.7756 0.7562 0.7547 0.7682 0.7497 0.7473 0.7554
ZSL 0.7284 0.7159 0.7123 0.7230 0.7102 0.7078 0.7167
Multiclass Hope Speech Detection
Llama3 FSL 0.5508 0.5225 0.5134 0.4647 0.3999 0.3756 0.5225
ZSL 0.3928 0.3645 0.3551 0.3457 0.3004 0.2819 0.3674
GPT-4 FSL 0.4993 0.4712 0.4639 0.4214 0.3937 0.3866 0.4712
ZSL 0.4547 0.4285 0.4196 0.4001 0.3792 0.3657 0.4285
Fine‑tuning domain‑specific transformers consistently delivers the strongest and most balanced performance across both languages (Tables 6, 8). In the binary task, English RoBERTa reaches a macro $F_{1}$ of 86.5%, while its Spanish counterpart Albert attains 84.2%; the weighted–macro divergences are $\leq 0.3$% in either language, confirming uniform treatment of Hope and Not‑Hope. Scaling to five classes reduces every metric by roughly 10% for English (RoBERTa macro $F_{1}$=75.9%) and about 12% for Spanish (Albert macro $F_{1}$=72.6%), yet the weighted–macro gaps stay below 3%, indicating that the extra difficulty is spread fairly rather than concentrating on the minority labels Sarcasm, Unrealistic Hope, and Realistic Hope. Prompt‑based large‑language‑model baselines (Tables 7, 9) narrow the margin in the binary setting—GPT‑4 with ten in‑context demonstrations yields weighted $F_{1}$ scores of 77.9% (English) and 75.5% (Spanish) but still lag the best fine‑tuned transformer by 8–9% and exhibit macro scores 5–9% lower than their own weighted figures, signalling over‑reliance on the dominant Not‑Hope class. Zero‑shot prompting widens these disparities further, bringing five‑fold macro $F_{1}$ down to 36–40% in both languages. Overall, the empirical evidence shows that supervised transformer adaptation not only yields higher aggregate accuracy ($\approx$79English, $\approx$73Spanish, five‑fold) but also preserves class balance, whereas few‑ and zero‑shot prompting—while informative as a reasoning baseline—remains insufficient for the fine‑grained distinctions mandated by the PolyHope V2 dataset.
Several interacting factors explain the systematic gap observed in Tables 79. First, the transformers were discriminative models explicitly fine‑tuned on PolyHope tweets; gradient updates aligned their internal representations with the lexical idiosyncrasies of X/Twitter, the pragmatic cues that mark sarcasm, and the subtle distinctions among our three hope sub‑types. GPT‑4 and Llama-3, by contrast, were deployed in a zero‑ or few‑shot generative regime: without parameter adaptation they relied on priors learned from broad, largely formal corpora where the boundary between “realistic” and “unrealistic” hope is seldom annotated. Second, the demonstration pool was deliberately kept minimal (ten examples per class) to probe “out‑of‑the‑box” reasoning. This is dwarfed by the stylistic diversity and label noise of social‑media text, so the LLMs struggle to infer robust class‑conditional cues, particularly for the minority labels that constitute $<10$% of either language’s corpus.
Fine‑tuned Pre-trained Language Models (PLMs), in contrast, see every training instance and therefore learn token‑level features even for sparse categories. Third, prompt‑based classification is highly sensitive to surface form: small edits in wording or label order can shift accuracy by more than 5% [46]. Our study used a single handcrafted template to ensure comparability, but that template may sit at a sub‑optimal point in the LLM’s likelihood landscape. Supervised Pre-trained Language Models (PLMs) avoid this volatility by emitting logits directly over a fixed label vocabulary. Moreover, the generative output must be parsed back into a discrete class, introducing an additional failure channel for example, returning wishful thinking instead of Unrealistic Hope or attaching extraneous punctuation which the evaluation counts as error even when the underlying reasoning is partially correct.
Finally, sarcastic hope often relies on culture specific pragmatic signals that differ between English and Spanish; without fine‑tuning, the LLMs lack exposure to these bilingual nuances, reducing macro scores in both languages. Hence, the observed deficiencies do not necessarily imply limited reasoning capacity in GPT‑4 or Llama 3; rather, they reflect domain mismatch, extremely small in‑context support, prompt sensitivity, and generative‑to‑discriminative conversion overheads—constraints we accepted so that the LLM runs would serve as a clean baseline of unadapted reasoning on this complex multilingual taxonomy.
6 Discussion and Analysis
Figure 1 and 2 shows the multi-class confusion matrix between the best performing transfomer models. Figure 1 (English) and Figure 2 (Spanish) show that the best transformer already separates the two polar classes extremely well—Sarcasm is identified in more than 94% of cases in both languages, and Not‑Hope achieves recalls of 85.4% (3,448/4,081) in English and 85.8% (8,395/9,788) in Spanish—yet the model falters when deciding which variety of sincere hope is being expressed. In English almost one quarter of all Generalized Hope tweets (295/2,335, 12.6%) are downgraded to Not‑Hope, while the three hope sub‑types frequently trade places with one another (e.g., 20.9% of Realistic Hope is predicted as Generalized, 11.7% of Unrealistic Hope as Realistic). The Spanish matrix magnifies the same phenomenon: although Generalized Hope still attains a respectable 72.8% recall (3,646/5,007), the model misroutes nearly one fifth of those instances to Not‑Hope and, most strikingly, recognises fewer than half of the Realistic Hope tweets correctly (49.5%, 1,001/2,024), dispersing the remainder almost evenly across the other hope labels. These error patterns suggest that the network has learned robust cues for broad polarity and for pragmatic irony, but lacks the contextual or world‑knowledge signals required to gauge how grounded a hopeful statement is, a limitation that grows sharper under the stronger class imbalance of the Spanish data.
Figure 1: The figure shows the confusion matrix of the best-performing transformer model (RoBERTa) in English.
Figure 2: The figure shows the confusion matrix of the best-performing transformer model (RoBERTa) in Spanish.
To understand why the model stumbles on the three fine‑grained hope labels, we manually inspected 385 mis‑classified tweets from one English fold ($\approx 20\%$ of its 1,903 test items). Three recurrent patterns emerged:
  • Lexical “hope” triggers without genuine optimism. Many Not‑Hope posts contain aspirational verbs in a negative frame, e.g. We yearn for change, but nothing ever improves” (true Not‑Hope, predicted Generalized Hope). The vocabulary cue yearn” activates the hope prototype even though the surrounding discourse is pessimistic.
  • Implicit stance overrides explicit markers. Political commentary that affirms a group yet attacks its opponents is often annotated as Generalized Hope but judged by the model as Not‑Hope. Example: “#Woke people accuse conservatives of narrow‑mindedness; I still believe we can find common ground” (true Generalized Hope, predicted Not-Hope). The optimism is subtle, buried inside an adversarial sentence.
  • Boundary fuzziness between hope sub‑types. Tweets that express a concrete but weakly grounded desire swing between Generalized, Realistic, and Unrealistic Hope. For instance, Hey #USER, hope you’ll read this on air tomorrow!” (true Generalized Hope, predicted Realistic Hope) lacks external evidence and was annotated as Generalized, yet the model judged the request sufficiently specific to be Realistic. Conversely, progress‑update posts such as Still trying to finish Chapter 5 before July; wish me luck” (true Realistic Hope, predicted Generalized Hope) are grounded in an attainable goal but resemble generic motivational tweets in form.
These observations align with the quantitative error matrix: the classifier already recognises overt polarity and sarcastic cues, but its token‑level heuristics cannot reliably infer the speaker’s degree of realism. Future work should therefore incorporate temporal or evidential features—e.g. futurity markers, probability adverbs, or external knowledge about plausibility—to sharpen the distinctions among Generalized, Realistic, and Unrealistic Hope.
7 Conclusion
This paper introduces PolyHope V2, a novel bilingual dataset and modeling framework aimed at detecting fine-grained categories of hope speech—including sarcasm—across English and Spanish tweets. Our results demonstrate that fine-tuned transformer models consistently outperform zero-shot and few-shot large language models, particularly in maintaining class balance and accurately distinguishing among nuanced subtypes of hope. While LLMs like GPT-4 and Llama 3 show promise in binary classification tasks, their performance declines sharply in more granular multi-class settings due to their lack of domain adaptation and sensitivity to prompt design. The inclusion of sarcasm as a separate category further sharpens the challenge, revealing the limitations of current models in interpreting pragmatic cues. Our analysis underscores the need for richer annotations, cultural grounding, and external knowledge integration to advance the state of hope-speech detection. PolyHope V2 thus provides a strong empirical and methodological basis for future multilingual emotion detection research, bridging the gap between surface-level sentiment and deeper, context-aware emotional understanding.

### v2

@@ 3.1.2 Model Selection and Training Regimen
We conducted experiments on both English and Spanish datasets. The specific models used for each language are detailed in Tables 1 and 2. All models were fine-tuned using the AdamW optimizer with a learning rate of 3e-5 and a weight decay of 0.01. We trained for a maximum of 15 epochs per fold, employing early stopping with a patience of 3 epochs based on the performance on a held-out development set. The maximum sequence length was set to 100 tokens. A batch size of 32 was used for all experiments. The choice of these hyperparameters was guided by a coarse grid search on a separate development set, aiming to optimize performance while preventing overfitting.
Table 1: Pre-trained Transformer Models Used for English
@@
Table 2: Pre-trained Transformer Models Used for Spanish
@@
Through this fine-tuning exercise, we aim to establish strong, reproducible baselines for the PolyHope task. The results obtained will quantify the effectiveness of Transformer models for hope speech detection and, critically, sarcasm identification. This comparative analysis will provide a benchmark against which to evaluate the performance gains achieved by our proposed ZSL and FSL methodologies leveraging large language models.
@@ 3.2 Zero-Shot and Few-Shot Learning with Large Language Models
We selected GPT-4 [achiam2023gpt, bubeck2023sparks] and Llama 3 [grattafiori2024llama] for their superior performance in various natural language understanding and generation tasks. GPT-4, known for its advanced reasoning and generalization abilities, serves as a high-performing benchmark. Llama 3, a powerful open-source LLM, offers an alternative perspective and allows for greater control and customization.
@@ 3.2.2 Implementation Details
    messages=[{"role": "system", "content": "You are an NLP model trained for text classification."},
@@
By evaluating GPT-4 and Llama 3 in ZSL and FSL settings, we aim to assess the potential of LLMs for hope speech and sarcasm detection without task-specific training. The results will provide insights into the strengths and limitations of these models and inform future research directions in this area. We expect that FSL will outperform ZSL, and that GPT-4 and Llama 3 will exhibit different strengths and weaknesses due to their architectural differences and training data. A comparison with the fine-tuned Transformer baselines will reveal the trade-offs between these two approaches and highlight the potential for combining them to achieve even better performance.